The Bugatti Veyron Super Sport has a maximum speed of 268 miles per hour. The engine provides maximum forward force of 12,350 N on the 1900 kg car. Quadratic air resistance dominates, with a drag coecient of .86 kg/m. If the car starts from rest at one end of a 5 km long test track, can it reach a speed of 265 mph before it runs out of road? (Watch your units!).
Answer #1

fret - fengine - bu? = mudu 12350 - 0.86 02 = (1900) v dv udu - goo Jan 10 J 12350-03642 J 12350-0.8602 x-0 = 1900 en (12850-

v= 265 mph = 265x1609m 36008 = 118.44 mis 12350 - 0.86 (118.442) ? x = - 1104.65 lu 12350 x= 4160m = 4.16 km x < skom so befo